A Professional’s Guide to an Economical, Secure, and Functional Computing

Environment

Florin B. Manolache

                       florin@andrew.cmu.edu

Carnegie Mellon University Pittsburgh, Pennsylvania, USA

Overview

Computing Environment The High Qualification Model (HiQ) Structure & Management Open Source Software (OSS) Transition Cost Analysis Results !!!

Computing Environment Structure

RoEduNet Second International Conference 2003 – p.3/17

The High Qualification Model

PRINCIPLE: A computing environment is as good as the professionals supporting it directly.

The HiQ Directives

1. Expose the most qualified support professional.

The HiQ Directives

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

The HiQ Directives

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

The HiQ Directives

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

4. Users concerns should be addressed immediately.

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

4. Users concerns should be addressed immediately.

5. The software should be managed by professionals.

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

4. Users concerns should be addressed immediately.

5. The software should be managed by professionals.

6. Uniform hardware/software infrastructure.

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

4. Users concerns should be addressed immediately.

5. The software should be managed by professionals.

6. Uniform hardware/software infrastructure.

7. Refresh/rotate computers based on user needs.

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

4. Users concerns should be addressed immediately.

5. The software should be managed by professionals.

6. Uniform hardware/software infrastructure.

7. Refresh/rotate computers based on user needs.

8. No unplanned downtime is acceptable. (OSS)

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

4. Users concerns should be addressed immediately.

5. The software should be managed by professionals.

6. Uniform hardware/software infrastructure.

7. Refresh/rotate computers based on user needs.

8. No unplanned downtime is acceptable. (OSS)

9. R&D for professionals.

1. Expose the most qualified support professional.

2. Educate users by explaining problems and solutions.

3. No rigid arbitrary rules (what is "supported").

4. Users concerns should be addressed immediately.

5. The software should be managed by professionals.

6. Uniform hardware/software infrastructure.

7. Refresh/rotate computers based on user needs.

8. No unplanned downtime is acceptable. (OSS)

9. R&D for professionals.

10. No more consultants.

Implementation

First step: gradually start building the HiQ task force.

Implementation

First step: gradually start building the HiQ task force. Second step: start building custom software distributions matching the local needs and activities. (OSS)

Implementation

First step: gradually start building the HiQ task force.

Second step: start building custom software distributions matching the local needs and activities. (OSS)

Third step: start encouraging users to migrate to a standard hardware platform running the newly developed software distribution.

First step: gradually start building the HiQ task force.

Second step: start building custom software distributions matching the local needs and activities. (OSS)

Third step: start encouraging users to migrate to a standard hardware platform running the newly developed software distribution.

Fourth step: consolidate the environment.

First step: gradually start building the HiQ task force.

Second step: start building custom software distributions matching the local needs and activities. (OSS)

Third step: start encouraging users to migrate to a standard hardware platform running the newly developed software distribution.

Fourth step: consolidate the environment.

Fifth step: plan for the future.

Cost Analysis

The salary fund Software licensing costs Hardware costs User education costs Training of the support professionals Repair expenses Consulting expenses Server downtime Computer downtime Security expenses

Results

HiQ + OSS:

Great user satisfaction Lower costs by one order of magnitude per computer No downtime No security problems No data loss